Text Size:
Smaller Text Normal Text Larger Text

Facebook page

Twitter profile

YouTube channel

M1 and M2 aluminisation May 2002

Both mirrors were realuminised on the 27th of May 2002. The previous aluminisation took place in May 1997 for M1 and June 1995 for M2. The whole procedure started by dismounting the mirrors on the 24th of May, and the mirrors were transported back to the NOT on the 29th of May. While the mirrors were out, the telescope center section was machined to improve roundness.

Alignment was done on the 30th, and a report on the alignment procedure PDF file icon was produced by Anton Sørensen, CUO.

The throughput was tested with ALFOSC and StanCam on the 31st of May. Preliminary results for ALFOSC show the following sensitivity improvements:

  U_after / U_before = 1.59
  B                  = 1.51
  V                  = 1.36
  R                  = 1.32
  I                  = 1.29
Back to top Last modified: December 12 2022